# Zhaojian Yu, Xin Zhang, Ning Shang, Yangyu Huang, Can Xu, Yishujie Zhao, Wenxiang Hu, Qiufeng Yin Microsoft

### 20 Dec 2023 v1

@@ 1 Introduction
To enhance the performance of Code LLMs through instruction tuning, many existing methods of instruction data generation around the above two points. For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation. However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20]. Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs. Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
@@ 2 CodeOcean: Four-task Code-related Instruction Data
In this section, we the methodological details of our exploration. We firstly choose 4 representative coding tasks and collect raw code from open source dataset. For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning. The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
@@ 2.1 Tasks Setting
Code summarization (code-to-text). Code summarization task aims to summarize the given code in a brief utterance. Therefore, we take the raw code as input and reformulate the response of GPT-3.5-turbo in instruction format.
@@ 3 Training
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using 8 NVIDIA A100-80GB GPUs. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. The fine-tuning prompt is as follows:
@@ 4.2.1 Evaluation on Code Generation Task
Deepseek 6.7B - 49.4 60.6
WaveCoder-DS-6.7B [Uncaptioned image] 6.7B Deepseek 20K 62.8 (+13.4) 62.4 (+1.8)
@@
Deepseek-6.7B 29.9 29.2 39.0 29.2 25.0 21.9 29.0
Deepseek-CodeAlpaca-6.7B 49.4 51.8 45.1 48.8 44.5 31.7 45.2

### 26 Dec 2023 v2

@@ 1 Introduction
To enhance the performance of Code LLMs through instruction tuning, many existing methods for generating instructional data have been devised, focusing primarily on the two aforementioned aspects. For example, self-instruct [12], evol-instruct [13] employed the zero/few-shot ability of teacher LLM to generated instruction data, which present an amazing way for instruction data generation. However, these generation ways are too dependent on the performance of the teacher LLM and sometimes produces a large amount of duplicate data which will reduce the efficiency of fine-tuning [18, 19, 20]. Octopack [21] constructs a code instruction dataset leveraging the natural structure of Git commits to break away from dependence on teacher LLMs. Howerver, Ensuring the quality of data in git messages presents a considerable challenge, and the comprehensive screening of data through artificial filtering rules is often a complex task.
@@ 2 CodeOcean: Four-task Code-related Instruction Data
In this section, we introduce the methodological details of our exploration. We firstly choose 4 representative coding tasks and collect raw code from open source dataset. For each task, we employ GPT-3.5-turbo to generate instruction data for fine-tuning. The generation prompt is shown in Table 2 and the data generation process is shown in Figure 3.
@@ 3 Training
To obtain WaveCoder models, We choose StarCoder-15B, CodeLLaMa (7B and 13B), DeepseekCoder-6.7B as the base model and fine-tune all the base model for 3 epochs using NVIDIA A100-80GB GPU. For StarCoder-15B and CodeLLaMa-13B, we set the global batch size to 128 using Tensor Parallel and set the initial learning rate at 2e-5. For DeepseekCoder-6.7B and CodeLLaMa-7B, we set the global batch size to 256 using the Fully Sharded Data Parallel (FSDP) module from Pytorch and set the initial learning rate at 5e-5. The sequence length of our each instruction sample is less than 1024. The fine-tuning prompt is as follows:
@@ 4.2.1 Evaluation on Code Generation Task
DeepseekCoder 6.7B - 49.4 60.6
WaveCoder-DS-6.7B [Uncaptioned image] 6.7B DeepseekCoder 20K 62.8 (+13.4) 62.4 (+1.8)
@@
DeepseekCoder-6.7B 29.9 29.2 39.0 29.2 25.0 21.9 29.0
DeepseekCoder-CodeAlpaca-6.7B 49.4 51.8 45.1 48.8 44.5 31.7 45.2